# KAG: Boosting LLMs in Professional Domains via Knowledge Augmented Generation

### v1

@@ Abstract
The recently developed retrieval-augmented generation (RAG) technology enables the efficient construction of domain-specific applications. However, it faces limitations due to fuzzy retrieval processes, the "hallucination" problem of understanding and reasoning capabilities of general language models, and cascading losses in complex systems. These challenges hinder the effectiveness of specialized knowledge services. However, in scenarios such as scientific computing, medicine, and law, the accuracy of knowledge, the completeness of information, and the logical rigor of rules, time, and values are particularly critical. We Introduce professional domain knowledge service framework: Knowledge Augmented Generation(KAG) to improve generation and reasoning performance by bidirectionally enhancing large language model(LLM)s and knowledge graph(KG)s, including five key enhancements: 1) LLM-friendly knowledge semantic representation, 2) mutual indexing between knowledge graph and original chunks, 3) logical-form-guided hybrid reasoning and solving, 4) Knowledge alignment based on semantic reasoning, 5) Model for KAG. We compared KAG with existing RAG methods in multi-hop question answering. The results show that KAG performs significantly better than the state-of-the-art methods, with a relative improvement from 19.6% to 33.4% in F1. We apply KAG to two professional knowledge Q&A tasks of Ant Group, including E-Goverment Q&A and E-Health Q&A, and has achieved significant improvement in professionalism compared with NaiveRAG. We will soon natively support KAG on the open source KG engine OpenSPG, allowing developers to more easily build rigorous knowledge decision-making or convenient information retrieval services.
@@ 1 Introduction
Recently, the rapidly developing Retrieval-Augmented Generation (RAG)[1, 2, 3, 4, 5] technology, which utilizes external retrieval systems, has effectively improved the timeliness of large language models in acquiring task information and reduced the hallucination of answers. This allows for the efficient construction of applications in specific domains. However, general RAG is still constrained by the ambiguity of retrieval, the "hallucination" problem of the understanding and reasoning capabilities of general language models, and the cascading losses of complex systems. These issues hinder the professionalism of knowledge services, especially in fields such as law, healthcare, and scientific computing. To address these problems, Knowledge Graph, as a precise knowledge reasoning technology, are increasingly being integrated into the RAG framework. HippoRAG[6], DALK[7], and ToG 2.0[8] innovatively use triple-based inverted index to replace term-based inverted index, improving the indexing structure of documents, and making full use of graph computing related technologies to support document retrieval or summary aggregation.
Although these works organize documents using graph structures, allowing the system to perform better in multi-hop reasoning tasks, they still do not fully apply the specialized knowledge management capabilities of KGs to the system. This is reflected in the following aspects:
  • Compared with the knowledge graph constructed by domain experts, the automatically constructed structured knowledge only contains fragmented information from the document and lacks domain knowledge that is not mentioned in the documents. This missing domain knowledge serves as an intrinsic connection between document information and user queries, affecting the accuracy of problem-solving. For example, "industrial insurance", "unemployment insurance", and "housing provident fund" all fall under the category of "five insurances and one fund", but they might not be explicitly mentioned in the e-government service documentation. Consequently, when querying "five insurances and one fund", it may fail to retrieve the correct answer.
  • The graph structure constructed through OpenIE does not explicitly obtain classification, context and other information, which introduces a lot of noise, including polysemy, synonyms, differences in word granularity, sparse relationships, etc., which brings great challenges to both reasoning and retrieval, and also reduces the accuracy of retrieval.
  • The professional domain is highly sensitive to rules, numerical values, time, coordinates, numerical logic, or causal logic. For example, when analyzing data from sources like corporate financial reports, medical test results, or case descriptions to generate analytical messages or answer queries, it is necessary to capture one or a set of key metrics. This involves determining whether these metrics are showing an increasing or decreasing trend in chronological or distance order, and whether the magnitude of these changes exceeds normal fluctuation ranges. It is also important to identify potential causes of abnormal indicators and their possible impacts. These logical connections are intricately linked, and a small error can lead to significant misinterpretations. Therefore, there is a need for a mechanism or systematic language for the representation, retrieval, and controllable generation of professional logic.
  • The processes of index construction, inference retrieval, and answer generation pose challenges to the capabilities of the LLM models upon which the system relies.
To address the above challenges and meet the requirements of professional domain knowledge services, we hope to fully combine the complementary features of RAG and KG technologies, and then simultaneously utilize the understanding and generation capabilities of LLMs to improve the professional level of vertical domain knowledge services. In this paper, we propose a professional domain knowledge service framework: Knowledge Augmented Generation(KAG) to improve generation and reasoning performance by bidirectionally enhancing large language LLM and KG. As shown in Figure 1. We summarize our main contributions as follows:
@@
  • We proposed a LLM friendly knowledge representation framework LLMFriSPG. We refer to the hierarchical structure of data, information, and knowledge of DIKW to upgrade SPG to be friendly to LLMs, named LLMFriSPG, to make it compatible with schema-free information extraction and schema-constrained expert knowledge construction on the same knowledge type (such as entity type, event type), and supports the mutual indexing representation between graph structure and original text chunks, which facilitates the construction of graph-structure-based inverted index and facilitates the unified representation, reasoning, and retrieval of logical form.
@@
  • We proposed a logical-form-guided hybrid solving and reasoning engine. It includes two types of operators: reasoning and retrieval, transforming natural language questions into a problem-solving process that combines language and symbols. Each step in the process can utilize different operators such as exact match retrieval, text retrieval, numerical computation, or semantic reasoning, thereby achieving the integration of four distinct problem-solving processes: retrieval, KG reasoning, language reasoning, and numerical computation.
@@ 2 Approach
In this section, we will first introduce the overall framework of KAG, and then discuss five key technology in sections 2.1 to 2.5.
As shown in Figure 1, the KAG framework consists of three parts: KAG-Builder, KAG-Solver, and KAG-Model. The KAG-Builder is designed for building offline indexes; in this module, we proposed LLM Friendly Knowledge Representation framework and mutual indexing between knowledge structure and text chunk. In the module KAG-Solver we introduced a Logical-form-guided hybrid reasoning solver that integrates LLM reasoning, knowledge reasoning, and mathematical logic reasoning. Additionally, knowledge alignment by semantic reasoning is used to enhance the accuracy of knowledge representation and retrieval in both KAG-Builder and KAG-Solver. The KAG-Model optimizes the capabilities needed by each module based on a general language model, thereby improving the performance of all modules.
@@ 2.1 LLM Friendly Knowledge Representation
Figure 2: LLMFriSPG:A knowledge representation framework that is friendly to LLMs. Instances and concepts are separated to achieve more effective alignment with LLMs through concepts. In this paper, entity instances and event instances are collectively referred to as instances unless otherwise specified. SPG properties are divided into Knowledge and Information areas, also called static and dynamic area, which are compatible with decision-making expertise with strong schema constraints and document retrieval index knowledge with open information representation. The red dotted line represents the fusion and mining process from information to knowledge. The enhanced document chunk representation provides traceable and interpretable text context for LLMs.
@@
where $\mathcal{M}$ represents all types defined in LLMFriSPG, $\mathcal{T}$ represents all EntityType(e.g., Person in Figure 2), EventType classes and all pre-defined properties that are compatible with LPG syntax declarations. $\mathcal{C}$ represents all ConceptType classes, concepts and concept relations, it is worth noting that the root node of each concept tree is a ConceptType class that is compatible with LPG syntax(e.g., TaxoOfPerson in Figure 2), each concept node has a unique ConceptType class. $\mathcal{\rho}$ represents the inductive relations from instances to conecepts. $\mathcal{L}$ represents all executable rules defined on logical relations and logical concepts. For $\mathcal{\forall}{t}\in\mathcal{T}$:
@@
As is show in Figure 2, $p_{t}$ represents all properties and relations of type $t$, and $p_{t}^{c}$ represents the domain experts pre-defined part, $p_{t}^{f}$ represents the part added in an ad-hoc manner, $p_{t}^{b}$ represents the system built-in properties, such as supporting_chunks, descripiton, summary and belongTo. For any instance $e_{i}$, denote $typeof(e_{i})$ as $t_{k}$, and supporting_chunks represents the set of all text chunks containing instance ${e_{i}}$, the user defines the chunk generation strategy and the maximum length of the chunk in KAG builder phase, description represents the general descriptive information specific to class ${t_{k}}$. It is worth noting that the meaning of description added to the type ${t_{k}}$ and the instance ${e_{i}}$ is different, when description is attached to ${t_{k}}$, it signifies the global description for that type. Conversely, when it is associated with an instance ${e_{i}}$, it represents the general descriptive information for ${e_{i}}$ consistent with the orignal document context, description can effectively assist LLM in understanding the precise meaning of a specific instance or type, and can be used in tasks such as information extraction, entity linking, and summary generation. summary represents the summary of ${e_{i}}$ or ${r_{j}}$ in the original document context. belongTo represents the inductive semantics from instance to concept. Each EntityType or EventType can be associated with a ConceptType through belongTo. It is worth noting that, 1) $\mathcal{T}$ and $\mathcal{C}$ have different functions. The statement ${t}$ adopts the object-oriented principle to better match the representation of the LPG[12], and $\mathcal{C}$ is managed by a text-based concept tree. This article will not introduce the SPG semantics in detail. 2) $p_{t}^{c}$ and $p_{t}^{f}$ can be instantiated separately. That is, they share the same class declaration, but in the instance storage space, pre-defined static properties and realtime-added dynamic properties can coexist, and we also support instantiating only one of them. This approach can better balance the application scenarios of professional decision-making and information retrieval. General information retrieval scenarios mainly instantiate dynamic properties, while professional decision-making application scenarios mainly instantiate static properties. Users can strike a balance between ease of use and professionalism based on business scenario requirements. 3) $p_{t}^{c}$ and $p_{t}^{f}$ share the same conceptual terminology. Concepts are general common sense knowledge that is independent of specific documents or instances. Different instances are linked to the same concept node to achieve the purpose of classifying the instances. We can achieve semantic alignment between LLM and instances through concept graphs, and concepts can also be used as navigation for knowledge retrieval. see section 2.4 and 2.3 for details.
@@
In order to more accurately define the hierarchical representation of information and knowledge, as shown in 3, we denote $KG_{cs}$ as knowledge layer, which represents the domain knowledge that complies with the domain schema constraints and has been summarized, integrated, and evaluated. denote $KG_{fr}$ as information layer, which represents the graph structure information such as entities and relations obtained through information extraction. denote $RC$ as raw chunks layer, which represents the original document chunks after semantic segmentation. the $KG_{cs}$ layer fully complies with the SPG semantic specification and supports knowledge construction and logical rule definition with strict schema constraints, SPG requires that domain knowledge must have predefined schema constraints. It has high knowledge accuracy and logical rigor. However, due to its heavy reliance on manual annotation, the labor cost of construction is relatively high and the information completeness is insufficient. $KG_{fr}$ shares the same EntityTypes, Eventtypes and Conceptual system with $KG_{cs}$, and provides effective information supplement for $KG_{cs}$. At the same time, the supporting_chunks, summary, and description edges built between $KG_{fr}$ and $RC$ form an inverted index based on graph structure, making $RC$ an effective original-text-context supplement for $KG_{fr}$ and with high information completeness. As is show in the right part of figure 3, in a specific domain application, $R(KG_{cs})$, $R(KG_{fr})$, and $R(RC)$ respectively represent their knowledge coverage in solving the target domain problems. If the application has higher requirements for knowledge accuracy and logic rigorousness, it is necessary to build more domain structured knowledge and consume more expert manpower to increase the coverage of $R(KG_{cs})$. On the contrary, if the application has higher requirements for retrieval efficiency and a certain degree of information loss or error tolerance, it is necessary to increase the coverage of $R(KG_{fr})$ to fully utilize KAG’s automated knowledge construction capabilities and reduce expert manpower consumption.
@@ 2.2 Mutual Indexing
As illustrated in Figure 4, KAG-Builder consists of three coherent processes: structured information acquisition, knowledge semantic alignment and graph storage writer. The main goals of this module include: 1) building a mutual index between the graph structure and the text chunk to add more descriptive context to the graph structure, 2) using the concept semantic graph to align different knowledge granularities to reduce noise and increase graph connectivity.
@@ 2.2.2 Information Extraction with More Descriptive Context
Given a dataset, we use fine-tuning-free LLM(such as GPT-3.5, DeepSeek, QWen, etc,.) or our fine-tuned model Hum to extract entities, events, concepts and relations to construct $KG_{fr}$, subsequently, construct the mutual index structure between $KG_{fr}$ and $RC$, enabling cross-document links through entities and relations. This process includes three phases. First, it extracts the entity set $E=\{e_{1},e_{2},e_{3},...\}$ chunk by chunk, second, extracts the event set ${EV}=\{{ev}_{1},{ev}_{2},{ev}_{3},...\}$ associated to all entities and iteratively extracts the relation set $R=\{r_{1},r_{2},r_{3},...\}$ between all entities in $E$, finally, completes all hypernym relations between the instance and its spgClass. To provide more convenience for the subsequent Knowledge Alignment phase, and overcome the problem of low discrimination of knowledge phrases such as Wikidata and ConceptNet, in the entity extraction phase, we use LLMs to generate built-in properties description, summary, semanticType, spgClass, descripitonOfSemanticType by default for each instance e at one time, as shown in Figure 2, we store them in the e instance storage according to the structure of e.description,e.summary, <e, belongTo, semanticType> and <e, hasClass, spgClass>.

### v2

@@ Abstract
The recently developed retrieval-augmented generation (RAG) technology has enabled the efficient construction of domain-specific applications. However, it also has limitations, including the gap between vector similarity and the relevance of knowledge reasoning, as well as insensitivity to knowledge logic, such as numerical values, temporal relations, expert rules, and others, which hinder the effectiveness of professional knowledge services. In this work, we introduce a professional domain knowledge service framework called Knowledge Augmented Generation (KAG). KAG is designed to address the aforementioned challenges with the motivation of making full use of the advantages of knowledge graph(KG) and vector retrieval, and to improve generation and reasoning performance by bidirectionally enhancing large language models (LLMs) and KGs through five key aspects: (1) LLM-friendly knowledge representation, (2) mutual-indexing between knowledge graphs and original chunks, (3) logical-form-guided hybrid reasoning engine, (4) knowledge alignment with semantic reasoning, and (5) model capability enhancement for KAG. We compared KAG with existing RAG methods in multi-hop question answering and found that it significantly outperforms state-of-the-art methods, achieving a relative improvement of 19.6% on 2wiki and 33.5% on hotpotQA in terms of F1 score. We have successfully applied KAG to two professional knowledge Q&A tasks of Ant Group, including E-Government Q&A and E-Health Q&A, achieving significant improvement in professionalism compared to RAG methods. Furthermore, we will soon natively support KAG on the open-source KG engine OpenSPG, allowing developers to more easily build rigorous knowledge decision-making or convenient information retrieval services. This will facilitate the localized development of KAG, enabling developers to build domain knowledge services with higher accuracy and efficiency.
@@ 1 Introduction
Recently, the rapidly developing Retrieval-Augmented Generation (RAG)[1, 2, 3, 4, 5] technology, which utilizes external retrieval systems, has effectively improved the timeliness of large language models in acquiring task information and reduced the hallucination of answers. This allows for the efficient construction of applications in specific domains. In order to enhance the performance of the RAG system in multi-hop and cross-paragraph tasks, knowledge graph, renowned for high precision and strong logicality, have been introduced into the RAG technical framework, including GraphRAG[6], DALK[7], SUGRE[8], ToG 2.0[9], and HippoRAG[10].
Although RAG and its optimization have solved most of the hallucination problems caused by knowledge deficiency and updates, the generated text still lacks coherence and logic, rendering it incapable of producing correct and valuable answers, particularly in specialized domains such as law, medicine, and science where analytical reasoning is crucial. This shortcoming can be attributed to three primary reasons. Firstly, real-world business processes typically necessitate inferential reasoning based on the specific relationships between pieces of knowledge to gather information pertinent to answering a question. RAG, however, commonly relies on the similarity of text or vectors for retrieving reference information, which may lead to incomplete and repeated search results. secondly, real-world processes often involve logical or numerical reasoning, such as determining whether a set of data increases or decreases in a time series, and the next token prediction mechanism used by language models is still somewhat weak in handling such problems.
To address the above challenges and meet the requirements of professional domain knowledge services, we propose Knowledge Augmented Generation(KAG), which fully leverages the complementary characteristics of KG and RAG techniques. More than merely integrating graph structures into the knowledge base process, it incorporates the semantic types and relationships of knowledge graph and the commonly used Logical Forms from KGQA (Knowledge Graph Question Answering) into the retrieval and generation process. As shown in Figure 1, this framework involves the optimization of the following five modules:
@@
  • We proposed a LLM friendly knowledge representation framework LLMFriSPG. We refer to the hierarchical structure of data, information, and knowledge of DIKW to upgrade SPG to be friendly to LLMs, named LLMFriSPG, to make it compatible with schema-free information extraction and schema-constrained expert knowledge construction on the same knowledge type (such as entity type, event type), and supports the mutual-indexing representation between graph structure and original text chunks, which facilitates the construction of graph-structure-based inverted index and facilitates the unified representation, reasoning, and retrieval of logical form.
@@
  • We proposed a logical-form-guided hybrid solving and reasoning engine. It includes three types of operators: planning, reasoning and retrieval, transforming natural language questions into a problem-solving process that combines language and symbols. Each step in the process can utilize different operators such as exact match retrieval, text retrieval, numerical computation, or semantic reasoning, thereby achieving the integration of four distinct problem-solving processes: retrieval, KG reasoning, language reasoning, and numerical computation.
@@ 2 Approach
In this section, we will first introduce the overall framework of KAG, and then discuss five key enhancements in sections 2.1 to 2.5. As shown in Figure 1, the KAG framework consists of three parts: KAG-Builder, KAG-Solver, and KAG-Model. The KAG-Builder is designed for building offline indexes, in this module, we proposed a LLM Friendly Knowledge Representation framework and mutual-indexing between knowledge structure and text chunk. In the module KAG-Solver we introduced a Logical-form-guided hybrid reasoning solver that integrates LLM reasoning, knowledge reasoning, and mathematical logic reasoning. Additionally, knowledge alignment by semantic reasoning is used to enhance the accuracy of knowledge representation and retrieval in both KAG-Builder and KAG-Solver. The KAG-Model optimizes the capabilities needed by each module based on a general language model, thereby improving the performance of all modules.
@@ 2.1 LLM Friendly Knowledge Representation
Figure 2: LLMFriSPG:A knowledge representation framework that is friendly to LLMs. Instances and concepts are separated to achieve more effective alignment with LLMs through concepts. In this study, entity instances and event instances are collectively referred to as instances unless otherwise specified. SPG properties are divided into knowledge and information areas, also called static and dynamic area, which are compatible with decision-making expertise with strong schema constraints and document retrieval index knowledge with open information representation. The red dotted line represents the fusion and mining process from information to knowledge. The enhanced document chunk representation provides traceable and interpretable text context for LLMs.
@@
where, $\mathcal{M}$ represents all types defined in LLMFriSPG, $\mathcal{T}$ represents all EntityType(e.g., Person in Figure 2), EventType classes and all pre-defined properties that are compatible with LPG syntax declarations. $\mathcal{C}$ represents all ConceptType classes, concepts and concept relations, it is worth noting that the root node of each concept tree is a ConceptType class that is compatible with LPG syntax(e.g., TaxoOfPerson in Figure 2.), each concept node has a unique ConceptType class. $\mathcal{\rho}$ represents the inductive relations from instances to conecepts. $\mathcal{L}$ represents all executable rules defined on logical relations and logical concepts. For $\mathcal{\forall}{t}\in\mathcal{T}$:
@@
As is show in Figure 2, where, $p_{t}$ represents all properties and relations of type $t$, and $p_{t}^{c}$ represents the domain experts pre-defined part, $p_{t}^{f}$ represents the part added in an ad-hoc manner, $p_{t}^{b}$ represents the system built-in properties, such as supporting_chunks, descripiton, summary and belongTo. For any instance $e_{i}$, denote $typeof(e_{i})$ as $t_{k}$, and supporting_chunks represents the set of all text chunks containing instance ${e_{i}}$, the user defines the chunk generation strategy and the maximum length of the chunk in KAG builder phase, description represents the general descriptive information specific to class ${t_{k}}$. It is worth noting that the meaning of description added to the type ${t_{k}}$ and the instance ${e_{i}}$ is different, when description is attached to ${t_{k}}$, it signifies the global description for that type. Conversely, when it is associated with an instance ${e_{i}}$, it represents the general descriptive information for ${e_{i}}$ consistent with the orignal document context, description can effectively assist LLM in understanding the precise meaning of a specific instance or type, and can be used in tasks such as information extraction, entity linking, and summary generation. summary represents the summary of ${e_{i}}$ or ${r_{j}}$ in the original document context. belongTo represents the inductive semantics from instance to concept. Each EntityType or EventType can be associated with a ConceptType through belongTo. It is worth noting that, 1) $\mathcal{T}$ and $\mathcal{C}$ have different functions. The statement ${t}$ adopts the object-oriented principle to better match the representation of the LPG[14], and $\mathcal{C}$ is managed by a text-based concept tree. This article will not introduce the SPG semantics in detail. 2) $p_{t}^{c}$ and $p_{t}^{f}$ can be instantiated separately. That is, they share the same class declaration, but in the instance storage space, pre-defined static properties and realtime-added dynamic properties can coexist, and we also support instantiating only one of them. This approach can better balance the application scenarios of professional decision-making and information retrieval. General information retrieval scenarios mainly instantiate dynamic properties, while professional decision-making application scenarios mainly instantiate static properties. Users can strike a balance between ease of use and professionalism based on business scenario requirements. 3) $p_{t}^{c}$ and $p_{t}^{f}$ share the same conceptual terminology. Concepts are general common sense knowledge that is independent of specific documents or instances. Different instances are linked to the same concept node to achieve the purpose of classifying the instances. We can achieve semantic alignment between LLM and instances through concept graphs, and concepts can also be used as navigation for knowledge retrieval. the details are shown in section 2.4 and 2.3.
@@
In order to more accurately define the hierarchical representation of information and knowledge, as shown in 3, we denote $KG_{cs}$ as knowledge layer, which represents the domain knowledge that complies with the domain schema constraints and has been summarized, integrated, and evaluated. denote $KG_{fr}$ as graph information layer, which represents the graph data such as entities and relations obtained through information extraction. denote $RC$ as raw chunks layer, which represents the original document chunks after semantic segmentation. the $KG_{cs}$ layer fully complies with the SPG semantic specification and supports knowledge construction and logical rule definition with strict schema constraints, SPG requires that domain knowledge must have pre-defined schema constraints. It has high knowledge accuracy and logical rigor. However, due to its heavy reliance on manual annotation, the labor cost of construction is relatively high and the information completeness is insufficient. $KG_{fr}$ shares the same EntityTypes, Eventtypes and Conceptual system with $KG_{cs}$, and provides effective information supplement for $KG_{cs}$. Meanwhile, the supporting_chunks, summary, and description edges built between $KG_{fr}$ and $RC$ form an inverted index based on graph structure, making $RC$ an effective original-text-context supplement for $KG_{fr}$ and with high information completeness. As is show in the right part of figure 3, in a specific domain application, $R(KG_{cs})$, $R(KG_{fr})$, and $R(RC)$ respectively represent their knowledge coverage in solving the target domain problems. If the application has higher requirements for knowledge accuracy and logic rigorousness, it is necessary to build more domain structured knowledge and consume more expert manpower to increase the coverage of $R(KG_{cs})$. On the contrary, if the application has higher requirements for retrieval efficiency and a certain degree of information loss or error tolerance, it is necessary to increase the coverage of $R(KG_{fr})$ to fully utilize KAG’s automated knowledge construction capabilities and reduce expert manpower consumption.
@@ 2.2 Mutual Indexing
As illustrated in Figure 4, KAG-Builder consists of three coherent processes: structured information acquisition, knowledge semantic alignment and graph storage writer. The main goals of this module include: 1) building a mutual-indexing between the graph structure and the text chunk to add more descriptive context to the graph structure, 2) using the concept semantic graph to align different knowledge granularities to reduce noise and increase graph connectivity.
@@ 2.2.2 Information Extraction with More Descriptive Context
Given a dataset, we use fine-tuning-free LLM(such as GPT-3.5, DeepSeek, QWen, etc,.) or our fine-tuned model Hum to extract entities, events, concepts and relations to construct $KG_{fr}$, subsequently, construct the mutual-indexing structure between $KG_{fr}$ and $RC$, enabling cross-document links through entities and relations. This process includes three steps. First, it extracts the entity set $E=\{e_{1},e_{2},e_{3},...\}$ chunk by chunk, second, extracts the event set ${EV}=\{{ev}_{1},{ev}_{2},{ev}_{3},...\}$ associated to all entities and iteratively extracts the relation set $R=\{r_{1},r_{2},r_{3},...\}$ between all entities in $E$, finally, completes all hypernym relations between the instance and its spgClass. To provide more convenience for the subsequent Knowledge Alignment phase, and overcome the problem of low discrimination of knowledge phrases such as Wikidata[15] and ConceptNet[16], in the entity extraction phase, we use LLMs to generate built-in properties description, summary, semanticType, spgClass, descripitonOfSemanticType by default for each instance e at one time, as shown in Figure 2, we store them in the e instance storage according to the structure of e.description,e.summary, <e, belongTo, semanticType> and <e, hasClass, spgClass>.